# Natural Language Processing for the Legal Domain: A Survey of Tasks, Datasets, Models, and Challenges

### v1

@@ Abstract.
Natural Language Processing is revolutionizing the way legal professionals and laypersons operate in the legal field. The considerable potential for Natural Language Processing in the legal sector, especially in developing computational tools for various legal processes, has captured the interest of researchers for years. This survey follows the Preferred Reporting Items for Systematic Reviews and Meta-Analyses framework (Page et al., 2021), reviewing 148 studies, with a final selection of 127 after manual filtering. It explores foundational concepts related to Natural Language Processing in the legal domain, illustrating the unique aspects and challenges of processing legal texts, such as extensive document length, complex language, and limited open legal datasets. We provide an overview of Natural Language Processing tasks specific to legal text, such as Legal Document Summarization, legal Named Entity Recognition, Legal Question Answering, Legal Text Classification, and Legal Judgment Prediction. In the section on legal Language Models, we analyze both developed Language Models and approaches for adapting general Language Models to the legal domain. Additionally, we identify 15 Open Research Challenges, including bias in Artificial Intelligence applications, the need for more robust and interpretable models, and improving explainability to handle the complexities of legal language and reasoning.
copyright: acmlicensedjournal: CSURccs: General and reference Surveys and overviewsccs: Applied computing Lawccs: Computing methodologies Natural language processingccs: Computing methodologies Natural language generationccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Supervised learningccs: Computing methodologies Unsupervised learningccs: Computing methodologies Reinforcement learningccs: Computing methodologies Multi-task learningccs: Computing methodologies Machine learning approachesccs: Computing methodologies Artificial intelligenceccs: Information systems Language modelsccs: Information systems Retrieval tasks and goalsccs: Information systems Question answeringccs: Information systems Clustering and classificationccs: Information systems Summarization
@@ 1. Introduction
In recent years, advancements in Natural Language Processing (NLP) have significantly impacted the legal domain by simplifying complex tasks such as Legal Document Summarization (LDS), enhancing legal text comprehension for laypersons, and improving Legal Question Answering (LQA) and Legal Judgment Prediction (LJP) (Savelka et al., 2023; Pont et al., 2023; Janatian et al., 2023; Gesnouin et al., 2024; Chlapanis et al., 2024; Jiang et al., 2024; Li et al., 2022a). These improvements are primarily attributed to advancements in neural network architectures, such as transformer models (Vaswani et al., 2017). NLP techniques now enable machines to generate text, answer legal questions, drafting a regulation, and simulate legal reasoning, which can revolutionize legal practices (Janatian et al., 2023). Applications like contract review (Graham et al., 2023; Mamooler et al., 2022; Maroudas et al., 2022; Tuggener et al., 2020) and case prediction (Niklaus et al., 2021; Vianna et al., 2023) have been automated to a large extent, speeding up processes, reducing human error, and cutting operational costs. Additionally, the use of NLP allows lawyers and legal professionals to reduce their workload, enhance efficiency, and minimize errors in decision-making processes (Savelka et al., 2023). Despite the rapid development of NLP, challenges remain in processing lengthy documents, understanding complex language, and navigating complicated document structures (Yuen Iu and Man-Yi Wong, 2023; Vianna et al., 2023; Sovrano et al., 2021; Nguyen et al., 2024; Jiang et al., 2024; Ganguly et al., 2023; Bhattacharya et al., 2019), yet the promise of Large Language Models (LLMs) enhances the efficiency, accessibility, and precision of legal services.
@@
Despite these advantages, the integration of NLP in the legal domain is not without challenges such as, biases and unfairness, and explainability issues. (Tamkin et al., 2021; Sheng et al., 2019; Deroy and Maity, 2023). The use of Artificial Intelligence (AI) in legal applications must follow strict standards of accuracy, fairness, and transparency, given the potential impact on clients’ lives and rights.
@@
This survey article explores the current landscape of NLP applications within the legal domain. It discusses its potential benefits and the practical challenges it poses. NLP is a broad field covering a wide range of techniques for processing, analyzing, and understanding human language. By examining the latest advancements and applications of NLP in law, this article provides a comprehensive overview of the field. Figure 1 summarizes the scope of the survey and categorise the research into several areas: LQA, LJP, Legal Text Classification (LTC), LDS, legal Named Entity Recognition (NER), legal corpora, and legal Language Models (LMs). Each category lists relevant projects and papers and shows the work being done in each sub-field. Notably, there is comparatively less research in NER and legal corpora, whereas LDS and LQA have seen extensive research activity, with a substantial number of datasets and research contributions. This summary provides an overview of how NLP techniques are applied to various challenges in the legal domain and offers insights into future directions of AI in legal practice.
{forest}
forked edges, for tree= draw, minimum size=1cm, semithick, rounded corners, align = center, inner sep = 3mm, edge = semithick, -stealth, grow = east, l sep = 12mm, fork sep=6mm, [Legal
NLP [LQA [Huang et al. (2020); Sovrano et al. (2020); Khazaeli et al. (2021)
Zhong et al. (2020c); Askari et al. (2022)
Zhang et al. (2023b); Louis et al. (2024)
Sovrano et al. (2024); Yuan et al. (2023); Askari et al. (2024)
Chen et al. (2023); Büttner and Habernal (2024); Sovrano et al. (2021) ] ] [LJP [Luo et al. (2017); Zhong et al. (2018); Ye et al. (2018)
Yang et al. (2019); Chalkidis et al. (2019a); Medvedeva et al. (2020)
Zhong et al. (2020a); Xu et al. (2020)
Ma et al. (2021); Almuslim and Inkpen (2022); Feng et al. (2022)
Tong et al. (2024); Zhang et al. (2023a); Liu et al. (2023)
Strickson and De La Iglesia (2020); Niklaus et al. (2021)
Semo et al. (2022) ] ] [LTC [Elnaggar et al. (2018); Wei et al. (2018); Lee and Lee (2019)
Bambroo and Awasthi (2021); Song et al. (2022)
Fragkogiannis et al. (2023); Wang et al. (2022); Mamooler et al. (2022)
Chalkidis et al. (2019b); Tuggener et al. (2020); Chalkidis et al. (2021)
Papaloukas et al. (2021); Graham et al. (2023); Nguyen et al. (2021) ] ] [LDS [Gelbart and Smith (1991a, b); Moens et al. (1999); Farzindar (2004)
Polsley et al. (2016); Zhong et al. (2019); Schraagen et al. (2022)
Zhong and Litman (2022); Moro et al. (2023); Jain et al. (2024)
Liu et al. (2024); Shen et al. (2022) ] ] [Legal NER [ Dozier et al. (2010); Păis et al. (2021); Smădu et al. (2022)
Leitner et al. (2020); Au et al. (2022); Kalamkar et al. (2022) ] ] [Legal corpora [ Zheng et al. (2021); Chalkidis et al. (2022a, b)
Xiao et al. (2018); Henderson et al. (2022); Rabelo et al. (2022)
Barale et al. (2023); Niklaus et al. (2023); Park and James (2023)
Goebel et al. (2024); Östling et al. (2024); Niklaus et al. (2024) ] ] [Legal LMs [Models [ Chalkidis et al. (2020); Xiao et al. (2021)
Colombo et al. (2024); Shi et al. (2024)
Al-qurishi et al. (2022) ] ] [Methods [ Li et al. (2022a); Mamakas et al. (2022) ] ] ] ]
Figure 1. An overview of the research areas in Legal NLP and the key publications in the survey.
[…]
Hierarchical diagram of existing research in Legal NLP.
To provide a comprehensive understanding of the existing research on integrating AI within the legal domain, we present an overview of recent literature reviews, as summarized in Table 1. Most survey papers on intelligent legal systems focus either on traditional NLP technologies for specific tasks like LJP and LDS or take a broader approach but still miss certain aspects. As illustrated in Table 1, there seems to be a gap in survey papers that thoroughly examine all facets of this multidisciplinary field. Our current work aims to bridge this gap by offering a comprehensive survey of all NLP tasks, existing datasets and corpora, and LMs in the legal domain. We use ✓ to indicate papers that study the most of the existing research on each subject in legal NLP. Papers that do not address a subject receive ✗, and those that partially cover specific subjects along with their datasets or LMs are marked with –.
Table 1. Comparison of existing surveys with this work, which shows the covered topics of each survey.
References Covered subjects Published year Description
Legal Dataset
NLP Tasks in Legal
Legal LM
Large Legal Corpus
Dias et al. (2022) 2022 In this work, researchers did not only concentrate on the legal domain. Instead, they elaborated on foundational concepts of AI and NLP. They briefly explored the applications of NLP within the legal field but did not delve deeply into legal datasets, specific NLP tasks in the legal domain, or Legal LLMs. In contrast, our work provides a comprehensive analysis of all NLP tasks within the legal domain, including LQA, LDS, LTC. Additionally, our study covers large legal corpora and thoroughly examines the datasets available for each legal NLP task, areas that were not fully addressed in this work.
Sun (2023) 2023 Sun explored a limited number of research projects in the field of LLMs and the legal domain. It focused on two key NLP tasks within this field: LJP and statutory reasoning. Additionally, it examined three datasets and two LLMs relevant to these domain. The difference is that our work provides a comprehensive overview of all NLP tasks in the legal domain, as well as large legal corpora and datasets for each task, which were not fully covered in Sun’s study.
Cui et al. (2023) 2023 This paper comprehensively reviewed on LJP task. Authors analyzed 43 LJP datasets in 9 different languages. They summarized 16 evaluation metrics to evaluate three NLP tasks (text classification, text generation, and text regression) in LJP. For LMs, they explored existing Pre-trained Language Models (PLMs). Unlike this paper, our work provides comprehensive coverage of all NLP tasks in the legal domain and includes an exploration of large legal corpora.
Anh et al. (2023) 2023 This survey give a short explanation regarding challenges in legal language and how LLMs try to overcome the challenges. Then, the authors summarized six NLP tasks in the legal domain that can be addressed by LLMs. In terms of ethical and legal considerations, they discussed ‘bias and fairness’, ‘privacy and confidentiality’, ‘intellectual property’, ‘explainability and transparency’, and ‘responsible use’. The difference is that our work focuses on existing methods for all Legal NLP tasks, including LJP and LDS, available datasets for each task, and large legal corpora for pre-training and fine-tuning, whereas they concentrated on ethical and legal considerations and the impact of LLMs on NLP in legal texts.
Ganguly et al. (2023) 2023 Ganguly and his colleagues presented a comprehensive review at ECIR 2023 and discuss several key areas including the processing challenges of legal text, such as NER and sentence boundary detection. They traced the historical evolution of AI and law research from the 1980s, highlighted recent developments in NLP and IR techniques with a focus on the architectures of PLMs, and conducted a detailed survey of current issues and advancements in legal IR and NLP tasks like LDS and LJP. The review also included perspectives from the industry. In contrast to their survey, we also explore LQA and LTC tasks, which they did not cover, and examine large legal corpora, an area they did not address.
Chen et al. (2024a) 2024 This survey study focuses on LLMs in three distinct fields: finance, healthcare, and law. Although it attempts to cover all aspects of the legal domain and LLMs, the broad scope of addressing three expansive topics has resulted in a less thorough examination of many specific research cases within the legal field. In addition, it did not cover large legal corpora which are used for pre-training and fine-tuning purposes. Also, we explore methods for improving the efficiency of LLMs in legal domain.
Krasadakis et al. (2024) 2024 This survey study focuses on challenges and advances in some NLP tasks, such as NER and Relation Extraction. Unlike this survey, our study covers all legal NLP tasks, along with their corresponding datasets and large legal corpora. Additionally, we examine existing LLMs tailored for the legal domain.
The main difference between this work and previous surveys is that this survey aims to provide a more general description of all aspects of NLP tasks in the legal domain, rather than focusing solely on specific applications. The main contributions of this survey are summarized as follows: (1) This article extends beyond previous surveys by examining a broad spectrum of studies, and applications of legal NLP. By discussing datasets and large corpora in 24 languages and exploring the popular legal LMs, this survey establishes itself as an important resource in the field of legal NLP; (2) The survey offers an in-depth look at the challenges of integrating NLP with legal applications, with detailed discussions on technical solutions that tackle these issues, thereby enhancing understanding and encouraging further research in this evolving field. (3) This survey also highlights the existing research gaps in legal NLP, identifying areas that require further exploration and development, and providing a road-map for future research efforts in the legal NLP domain.
@@
This document is organized as follows. In Section 2, Background and Foundational Concepts, we provide a detailed overview of legal language and the basic principles of NLP as they apply to the legal domain. In Section 3, we briefly explain the research methodology of this work and how we extracted the resources. In section 4, we explore various NLP tasks that are tailored for legal applications and show their unique requirements and the methodologies employed to address them. These specialized tasks leverage advanced NLP techniques to process, analyze, and extract meaningful information from legal texts, thereby facilitating more efficient and accurate legal research and decision-making. Additionally, we delve into the datasets available for training and evaluating legal NLP tasks, emphasizing their characteristics and the implications they have for model performance. Following this, in Section 5, we explore the development of LMs that has been specifically adapted for the legal field. Finally, in Section 6, we address the key challenges associated with deploying NLP technologies in legal settings, discussing both current issues and potential solutions. Since this survey contains many acronyms, Table 2 provides the list of acronyms and their meanings to make it easier to follow.
@@
Acronyms Meaning Acronyms Meaning
AMR Abstract Meaning Representation LJP Legal Judgment Prediction
AL Active Learning LQA Legal Question Answering
AI Artificial Intelligence LTC Legal Text Classification
BERT Bidirectional Encoder Representations from Transformers LSTM Long Short-Term Memory
Bi-GRU Bidirectional Gated Recurrent Units ML Machine Learning
Bi-LSTM Bidirectional LSTM MLM Masked Language Modeling
CanLII Canadian Legal Information Institute MGAT Multi-Graph Attention Network
CA Case Analysis ML-LJP Multi-Law aware LJP
CRF Conditional Random Field MLP Multi-Layer Perceptron
CNN Convolutional Neural Networks MPBFN Multi-Perspective Bi-Feedback Network
DL Deep Learning MTL Multi-task Learning
DAG Directed Acyclic Graph NER Named Entity Recognition
EDUs Elementary Discourse Units ORC Open Research Challenge
ECtHR European Court of Human Rights PEFT Parameter-Efficient Fine-Tuning
FSCS Federal Supreme Court of Switzerland PLM Pre-trained Language Model
GRU Gated Recurrent Unit PRISMA Preferred Reporting Items for Systematic Reviews and Meta-Analyses
GAT Graph Attention Networks QA Question Answering
GESAN Graph-Based Evidence Retrieval and Aggregation Network RNN Recurrent Neural Networks
IR Information Retrieval RL Reinforcement Learning
JEC-QA Judicial Examination of Chinese Question Answering RLHF Reinforcement Learning from Human Feedback
KG Knowledge Graph RAG Retrieval-Augmented Generation
KD Knowledge-Driven SLM Small Language Model
LLM Large Language Model SVM Support Vector Machine
LM Language Model TF-IDF Term Frequency Inverse Document Frequency
LDS Legal Document Summarization

### v2

@@ Abstract.
Natural Language Processing (NLP) is revolutionising the way legal professionals and laypersons operate in the legal field. The considerable potential for NLP in the legal sector, especially in developing computational tools for various legal processes, has captured the interest of researchers for years. This survey follows the Preferred Reporting Items for Systematic Reviews and Meta-Analyses framework, reviewing 154 studies, with a final selection of 133 after manual filtering. It explores foundational concepts related to NLP in the legal domain, illustrating the unique aspects and challenges of processing legal texts, such as extensive document length, complex language, and limited open legal datasets. We provide an overview of NLP tasks specific to legal text, such as Legal Document Summarisation, legal Named Entity Recognition, Legal Question Answering, Legal Argument Mining, Legal Text Classification, and Legal Judgement Prediction. In the section on legal Language Models (LMs), we analyse both developed LMs and approaches for adapting general LMs to the legal domain. Additionally, we identify 16 Open Research Challenges, including bias in Artificial Intelligence applications, the need for more robust and interpretable models, and improving explainability to handle the complexities of legal language and reasoning.
Natural Language Processing, Artificial Intelligence, Legal Domain, Law
copyright: acmlicensedjournal: CSURccs: General and reference Surveys and overviewsccs: Applied computing Lawccs: Computing methodologies Natural language processingccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
In recent years, advancements in Natural Language Processing (NLP) have significantly impacted the legal domain by simplifying complex tasks, such as Legal Document Summarisation (LDS), Legal Argument Mining (LAM) enhancing legal text comprehension for laypersons, and improving Legal Question Answering (LQA) and Legal Judgement Prediction (LJP) (Savelka et al., 2023; Pont et al., 2023; Janatian et al., 2023; Gesnouin et al., 2024; Chlapanis et al., 2024; Jiang et al., 2024; Li et al., 2022; Huang et al., 2024). These improvements are primarily attributed to advancements in Neural Network (NN) architectures, such as transformer models (Vaswani et al., 2017). NLP techniques now enable machines to generate text, answer legal questions, drafting a regulation, and simulate legal reasoning, which can revolutionise legal practices (Janatian et al., 2023). Applications, such as contract review (Graham et al., 2023; Mamooler et al., 2022; Maroudas et al., 2022; Tuggener et al., 2020) and case prediction (Niklaus et al., 2021; Vianna et al., 2023) have been automated to a large extent, speeding up processes, reducing human error, and cutting operational costs (Zhong et al., 2020b). Additionally, the use of NLP allows lawyers and legal professionals to reduce their workload, enhance efficiency, and minimise errors in decision-making processes (Savelka et al., 2023; Huang et al., 2024). Despite the rapid development of NLP, challenges remain in processing lengthy documents, understanding complex language, and navigating complicated document structures (Yuen Iu and Man-Yi Wong, 2023; Vianna et al., 2023; Sovrano et al., 2021; Nguyen et al., 2024; Jiang et al., 2024; Ganguly et al., 2023; Bhattacharya et al., 2019; Huang et al., 2024), yet the promise of Large Language Models (LLMs) enhances the efficiency, fairness, and precision of legal services (Huang et al., 2024; Fagan, 2024).
@@
Despite these advantages, the integration of NLP in the legal domain is not without challenges such as, biases and unfairness, and explainability issues (Tamkin et al., 2021; Sheng et al., 2019; Deroy and Maity, 2023). The use of Artificial Intelligence (AI) in legal applications must follow strict standards of accuracy, fairness, and transparency, given the potential impact on clients’ lives and rights.
@@
This survey article explores the current landscape of NLP applications within the legal domain. It discusses its potential benefits and the practical challenges it poses. NLP is a broad field covering a wide range of techniques for processing, analysing, and understanding human language. By examining the latest advancements and applications of NLP in law, this article provides a comprehensive overview of the field. Table 1 summarises the scope of the survey and categorise the research into several areas: LQA, LJP, Legal Text Classification (LTC), LDS, legal Named Entity Recognition (NER), LAM, legal corpora, and legal Language Models (LMs). Each category lists relevant projects and papers and shows the work being done in each sub-field. Notably, there is comparatively less research in NER and legal corpora, whereas LDS and LQA have seen extensive research activity, with a substantial number of datasets and research contributions. This summary provides an overview of how NLP techniques are applied to various challenges in the legal domain and offers insights into future directions of AI in legal practice.
Legal Natural Language Processing
Language Models Methods Li et al. (2022); Mamakas et al. (2022)
Models Chalkidis et al. (2020); Xiao et al. (2021); Colombo et al. (2024); Shi et al. (2024); Al-qurishi et al. (2022)
Datasets For Pre-training Niklaus et al. (2024); Henderson et al. (2022)
Benchmarks Zheng et al. (2021); Chalkidis et al. (2022a, b); Xiao et al. (2018); Rabelo et al. (2022); Barale et al. (2023); Niklaus et al. (2023); Park and James (2023); Goebel et al. (2024); Östling et al. (2024)
Tasks Argument Mining Poudyal et al. (2020); Palau and Moens (2009); Habernal et al. (2024); Santin et al. (2023)
Named Entity Recognition Dozier et al. (2010); Păis et al. (2021); Smădu et al. (2022); Leitner et al. (2020); Au et al. (2022); Kalamkar et al. (2022)
Document Summarisation Gelbart and Smith (1991a, b); Moens et al. (1999); Farzindar (2004); Polsley et al. (2016); Schraagen et al. (2022); Zhong and Litman (2022); Moro et al. (2023); Jain et al. (2024); Liu et al. (2024); Shen et al. (2022)
Text Classification Elnaggar et al. (2018); Wei et al. (2018); Lee and Lee (2019); Bambroo and Awasthi (2021); Song et al. (2022); Wang et al. (2022); Mamooler et al. (2022); Chalkidis et al. (2019b); Tuggener et al. (2020); Chalkidis et al. (2021); Papaloukas et al. (2021); Graham et al. (2023); Nguyen et al. (2021); Bhattacharya et al. (2023); Galassi et al. (2024); Grabmair et al. (2015)
Judgement Prediction Luo et al. (2017); Zhong et al. (2018); Ye et al. (2018); Yang et al. (2019); Chalkidis et al. (2019a); Medvedeva et al. (2020); Zhong et al. (2020a); Xu et al. (2020); Niklaus et al. (2021); Ma et al. (2021); Feng et al. (2022); Semo et al. (2022); Tong et al. (2024); Zhang et al. (2023a); Liu et al. (2023)
Question Answering Huang et al. (2020); Sovrano et al. (2020); Khazaeli et al. (2021); Zhong et al. (2020c); Askari et al. (2022); Zhang et al. (2023c); Louis et al. (2024); Sovrano et al. (2024); Yuan et al. (2023); Askari et al. (2024); Chen et al. (2023); Büttner and Habernal (2024); Sovrano et al. (2021)
Table 1. An overview of the research areas in legal NLP and the key publications in the survey.
@@
This document is organised as follows. Firstly, In Section 2, we discuss previous surveys in this multidisciplinary domain. In Section 3, we provide a detailed overview of legal language and the basic principles of NLP as they apply to the legal domain. In Section 4, we briefly explain the research methodology of this work and how we extracted the resources. In sections 5-11, we explore various NLP tasks that are tailored for legal applications and show their unique requirements and the methodologies employed to address them. Additionally, we study the datasets available for training and evaluating legal NLP tasks, emphasising their characteristics and the implications they have for model performance. Following this, in Section 12, we investigate the development of LMs that has been specifically adapted for the legal field. Finally, in Section 13, we address the key challenges associated with deploying NLP technologies in legal settings, discussing both current issues and potential solutions. Since this survey contains many acronyms, Table 2 provides the list of acronyms and their meanings to make it easier to follow.
@@
Acronyms Meaning
CJEU Court of Justice of the European Union
ECHR European Court of Human Rights
FSCS Federal Supreme Court of Switzerland
JEC-QA Judicial Examination of Chinese Question Answering
LAM Legal Argument Mining
LDS Legal Document Summarisation
LJP Legal Judgement Prediction
LQA Legal Question Answering
LTC Legal Text Classification
ML-LJP Multi-Law aware Legal Judgement Prediction
2. Related Work
Several studies have examined the use of NLP in the legal domain, each focusing on different aspects. To provide a comprehensive understanding of the existing research on integrating AI within the legal domain, we present an overview of recent literature reviews, as summarised in Table 3. Most survey papers on intelligent legal systems focus either on traditional NLP technologies for specific tasks, such as LJP and LDS or take a broader approach but still miss certain aspects. As illustrated in Table 3, there seems to be a gap in survey papers that thoroughly examine all facets of this multidisciplinary field. Our current work aims to bridge this gap by offering a comprehensive survey of all NLP tasks, existing datasets and corpora, and LMs in the legal domain. We use ✓ to indicate papers that study the most of the existing research on each subject in legal NLP. Papers that do not address a subject receive ✗, and those that partially cover specific subjects along with their datasets or LMs are marked with –.
Dias et al. (2022) discussed AI and NLP concepts and their applications in the legal domain. The study did not analyse legal datasets, specific NLP tasks, or legal LLMs. This work, in contrast, reviews NLP tasks in the legal domain, including LQA, LDS, and LTC, along with legal datasets and corpora. Sun (2023) examined two NLP tasks in the legal domain, such as LJP and statutory reasoning. The study reviewed three datasets and two LLMs related to these tasks. Unlike Sun’s work, this survey covers a wider range of NLP tasks and includes legal corpora and datasets. Cui et al. (2023) focused on LJP, reviewing 43 datasets in nine languages. The study evaluated classification, text generation, and regression tasks. It also discussed pre-trained LMs used for LJP. This work differs by covering multiple NLP tasks, legal corpora, and dataset availability.
Anh et al. (2023) explored challenges in legal language processing and how LLMs can address them. The study summarised six NLP tasks and discussed ethical concerns, including bias, privacy, and transparency. While these issues are relevant, this survey focuses on existing NLP methods, datasets, and legal corpora for pre-training and fine-tuning.
Ganguly et al. (2023) surveyed legal text processing challenges, such as NER and sentence boundary detection. The study reviewed historical developments in AI and law research, as well as recent NLP advancements. It covered LDS and LJP but did not examine LQA, LTC, or legal corpora. Our work covers all these areas.
Chen et al. (2024a) studied LLMs in finance, healthcare, and law. While attempting to provide a broad view of LLM applications in the legal domain, the study’s scope resulted in a less detailed review of specific NLP tasks and datasets. Additionally, it did not examine large legal corpora or pre-training methods, aspects that our survey addresses.
Krasadakis et al. (2024) focused on challenges and advancements in some NLP tasks, such as NER and Relation Extraction. Unlike our study, which reviews all NLP tasks alongside datasets and legal corpora, their work primarily investigated existing LLMs for the legal domain.
Table 3. Overview of existing surveys on NLP in the legal domain.
References Covered Subjects in the Legal Domain Published Year
Dataset NLP Tasks LM Large Corpora
Dias et al. (2022) 2022
Sun (2023) 2023
Cui et al. (2023) 2023
Anh et al. (2023) 2023
Ganguly et al. (2023) 2023
Chen et al. (2024a) 2024
Krasadakis et al. (2024) 2024
The main difference between this work and previous surveys is that this survey aims to provide a more general description of all aspects of NLP tasks in the legal domain, rather than focusing solely on specific applications. The main contributions of this survey are summarised as follows: (1) This article extends beyond previous surveys by examining a broad spectrum of studies, and applications of legal NLP. By discussing datasets and large corpora in 24 languages and exploring the popular legal LMs, this survey establishes itself as an important resource in the field of legal NLP. (2) The survey offers an in-depth look at the challenges of integrating NLP with legal applications, with detailed discussions on technical solutions that tackle these issues, thereby enhancing understanding and encouraging further research in this evolving field. (3) This survey also highlights the existing research gaps in legal NLP, identifying areas that require further exploration and development, and providing a road-map for future research efforts in the legal NLP domain.